Divide 210 in the ratio of 3:5:7
Found 2 solutions by MathLover1, math_tutor2020:
Answer by MathLover1(20849)
Answer by math_tutor2020(3817)


The ratio 3:5:7 scales up to 3x:5x:7x for some positive real number x.

The pieces 3x, 5x and 7x combine to 210
The term "divide" is a bit misleading (in my opinion) because effectively we're adding those pieces and not dividing them.
Ie we're saying 3x+5x+7x = 210 and not 3x/5x = 210 for instance.
I think a better term might be "we're splitting 210 into pieces of ratio 3:5:7".
Though "divide" isn't technically wrong of course.

Anyways, we have this equation
3x+5x+7x = 210

Let's solve for x
3x+5x+7x = 210
15x = 210
x = 210/15
x = 14

Then we can say
3x = 3*14 = 42
5x = 5*14 = 70
7x = 7*14 = 98

Answer: The number 210 splits into 42+70+98
Those three values are in the ratio 42:70:98 which fully reduces to 3:5:7 when you divide all parts by 14.